# Arg-LLaDA: Argument Summarization via Large Language Diffusion Models and Sufficiency-Aware Refinement

### v3

@@ 5 Results and Analysis
To evaluate the effectiveness of our proposed Arg-LLaDA framework, we revisit the three core research questions posed earlier.
Arg-LLaDA significantly outperforms both zero-shot and fine-tuned large language models (LLMs) in generating high-quality argument summaries.
As shown in Table 2), Arg-LLaDA achieves the highest ROUGE-1 ($0.525$), ROUGE-2 ($0.271$), and ROUGE-L ($0.453$) on the ArgKP dataset, outperforming both fine-tuned (Gemma, LLaDA) and zero-shot (Gemini 2.5, DeepSeek-R1) baselines. These results are consistent across the ASE dataset, where Arg-LLaDA also leads in ROUGE-L ($0.349$), BLEURT ($0.580$), and BERTScore ($0.741$), indicating strong lexical and semantic overlap with human-written references. Compared to its closest competitors—Gemini 2.5 and DeepSeek-R1—Arg-LLaDA shows modest but consistent improvements across all dimensions, particularly in BLEURT and BERTScore, which are more sensitive to semantic adequacy and contextual fidelity.
Human evaluation (Figure 2) further corroborates these findings. Arg-LLaDA receives the highest average scores across all three dimensions: Coverage ($4.7$), Faithfulness ($4.6$), and Conciseness ($4.4$). These results suggest that Arg-LLaDA not only captures a broader spectrum of argumentative content but also avoids common issues such as hallucination or redundancy. In contrast, while models like Gemini 2.5 ($4.5/4.4/4.2$) and DeepSeek-R1 ($4.4/4.3/4.1$) exhibit strong performance, they tend to generate longer outputs that introduce occasional repetition or irrelevant detail, resulting in slightly lower conciseness scores.
@@
As shown in Table LABEL:tab:human_eval, Arg-LLaDA outperforms all baselines in the Faithfulness ($4.6$) and Conciseness ($4.4$) dimensions. These improvements can be attributed to the model’s explicit mechanism for identifying and filtering spans that lack sufficient argumentative support before generation. Compared to its fine-tuned counterparts (e.g., Gemma, $4.3/4.2$; LLaDA, $4.2/4.1$), Arg-LLaDA demonstrates superior control over hallucinated or redundant content. This is particularly evident when contrasting with zero-shot models like Claude or Qwen 3, whose summaries frequently exhibit ungrounded generalizations or verbose phrasing, as reflected in lower average ratings for faithfulness and conciseness (both <4.0<4.0).
@@ 6 Conclusion
In this work, we introduced Arg-LLaDA, a novel large language diffusion framework for argument summarization that integrates sufficiency-guided remasking and iterative regeneration. Our method addresses key limitations of existing approaches by enabling controllable refinement of summaries through a flexible masking controller and sufficiency diagnosis module. Empirical results across two benchmark datasets demonstrate that Arg-LLaDA consistently outperforms state-of-the-art baselines. These findings highlight the potential of diffusion-based self-refinement in enhancing the factuality and structure of argumentative summaries.
@@
Despite the promising results of Arg-LLaDA, several limitations warrant discussion. First, although we trained annotators with backgrounds in argument mining and provided clear evaluation guidelines, human judgments remain inherently subjective. Variations in interpretation or attention to detail may affect the consistency of annotation, especially for nuanced aspects like sufficiency and redundancy. Second, while our datasets cover a wide range of debate topics, some of these topics may contain controversial or sensitive content. But such content is not the focus of this work.
Appendix A Dataset
Appendix B Instruction for Zero-shot Inference.

### v4

@@ 5 Results and Analysis
To evaluate the effectiveness of our proposed Arg‑LLaDA framework, we revisit the three core research questions posed earlier.
Arg‑LLaDA significantly outperforms both zero-shot and fine-tuned large language models (LLMs) in generating high-quality argument summaries.
As shown in Table 2), Arg‑LLaDA achieves the highest ROUGE-1 ($0.525$), ROUGE-2 ($0.271$), and ROUGE-L ($0.453$) on the ArgKP dataset, outperforming both fine-tuned (Gemma, LLaDA) and zero-shot (Gemini 2.5, DeepSeek-R1) baselines. These results are consistent across the ASE dataset, where Arg‑LLaDA also leads in ROUGE-L ($0.349$), BLEURT ($0.580$), and BERTScore ($0.741$), indicating strong lexical and semantic overlap with human-written references. Compared to its closest competitors—Gemini 2.5 and DeepSeek-R1—Arg‑LLaDA shows modest but consistent improvements across all dimensions, particularly in BLEURT and BERTScore, which are more sensitive to semantic adequacy and contextual fidelity.
Human evaluation (Figure 2) further corroborates these findings. Arg‑LLaDA receives the highest average scores across all three dimensions: Coverage ($4.7$), Faithfulness ($4.6$), and Conciseness ($4.4$). These results suggest that Arg‑LLaDA not only captures a broader spectrum of argumentative content but also avoids common issues such as hallucination or redundancy. In contrast, while models like Gemini 2.5 ($4.5/4.4/4.2$) and DeepSeek-R1 ($4.4/4.3/4.1$) exhibit strong performance, they tend to generate longer outputs that introduce occasional repetition or irrelevant detail, resulting in slightly lower conciseness scores.
@@
As shown in Figure 2, Arg‑LLaDA outperforms all baselines in the Faithfulness ($4.6$) and Conciseness ($4.4$) dimensions. These improvements can be attributed to the model’s explicit mechanism for identifying and filtering spans that lack sufficient argumentative support before generation. Compared to its fine-tuned counterparts (e.g., Gemma, $4.3/4.2$; LLaDA, $4.2/4.1$), Arg‑LLaDA demonstrates superior control over hallucinated or redundant content. This is particularly evident when contrasting with zero-shot models like Claude or Qwen 3, whose summaries frequently exhibit ungrounded generalizations or verbose phrasing, as reflected in lower average ratings for faithfulness and conciseness (both <4.0<4.0). Importantly, our human evaluations demonstrate not only absolute performance gains but also high reliability and alignment with automatic metrics. Across five annotators, Krippendorff’s $\alpha$ (krippendorff2011computing) reaches substantial agreement for all three dimensions (Coverage $0.69$, Faithfulness $0.74$, Conciseness $0.66$). Furthermore, Soft scores exhibit strong monotonic trends with human judgments, confirming that our automatic metrics meaningfully reflect annotator preferences.
@@ 6 Conclusion
In this work, we introduced Arg‑LLaDA, a novel large language diffusion framework for argument summarization that integrates sufficiency-guided remasking and iterative regeneration. Our method addresses key limitations of existing approaches by enabling controllable refinement of summaries through a flexible masking controller and sufficiency diagnosis module. Empirical results across two benchmark datasets demonstrate that Arg‑LLaDA consistently outperforms state-of-the-art baselines. These findings highlight the potential of diffusion-based self-refinement in enhancing the factuality and structure of argumentative summaries.
@@
Despite the promising results of Arg‑LLaDA, several limitations warrant discussion. First, although we trained annotators with backgrounds in argument mining and provided clear evaluation guidelines, human judgments remain inherently subjective. Variations in interpretation or attention to detail may affect the consistency of annotation, especially for nuanced aspects like sufficiency and redundancy. Second, while our datasets cover a wide range of debate topics, some of these topics may contain controversial or sensitive content. But such content is not the focus of this work.
Appendix A Instruction for Zero-shot Inference.
@@
Appendix B Human Evaluation Guidelines
To complement automatic metrics with more semantically grounded assessments, we employ a controlled human evaluation procedure following established practices in argument summarization. This appendix provides the full annotation guidelines, evaluator instructions, and quality-control measures used in our study.
B.1 Annotator Background and Training
All summaries were evaluated by three trained annotators with strong English proficiency, i.e. native English speakers or hold a Master’s degree from an English-speaking country.
Annotators first received a guideline document detailing the evaluation criteria and were asked to review example summaries with explanations. We then conducted a calibration round of 15 practice examples to harmonize interpretation of the rubric; disagreements were discussed and resolved before the official annotation began.
B.2 Evaluation Setup
For each instance, annotators were presented with:
  • the input claim,
  • the set of retrieved evidence spans, and
  • a randomly ordered, anonymized list of model-generated summaries
to prevent ordering bias and model-identification effects.
Annotators rated each summary independently using a 5-point Likert scale (1 = very poor, 5 = excellent) along three dimensions defined below.
B.3 Rating Rubric
Coverage (1–5 Scale)
  • 5 – Excellent: Captures all major arguments and viewpoints with no omissions.
  • 4 – Good: Covers most key points, with only minor or nuanced elements missing.
  • 3 – Fair: Conveys the main idea but omits several relevant viewpoints or supporting arguments.
  • 2 – Poor: Misses many important aspects; only captures a small portion of the argument space.
  • 1 – Very Poor: Fails to reflect the input’s arguments; largely incomplete or irrelevant.
Faithfulness (1–5 Scale)
  • 5 – Excellent: All statements are fully grounded in the input; no hallucinations or distortions.
  • 4 – Good: Mostly faithful with only minor unsupported phrasing or slight overgeneralization.
  • 3 – Fair: Generally accurate but contains noticeable paraphrasing errors or weakly supported claims.
  • 2 – Poor: Includes several unsupported, incorrect, or contradictory statements.
  • 1 – Very Poor: Largely unfaithful; contains substantial hallucinations or fabrications.
Conciseness (1–5 Scale)
  • 5 – Excellent: Succinct and clear; no redundancy; conveys the full argument with minimal wording.
  • 4 – Good: Mostly concise; contains only minor repetition of claim or slightly verbose phrasing.
  • 3 – Fair: Understandable but includes some redundancy or unnecessary detail.
  • 2 – Poor: Noticeably verbose, repetitive, or cluttered; key content is diluted.
  • 1 – Very Poor: Highly redundant or rambling; lacks clarity and wastes space.
Appendix C Detailed Experiment Result
GPT-4 GPT-o3 Gemini 2.5 Claude DeepSeek-R1 LLaMA 3.1 LLaDA Mistral $sP_{\text{BLEURT}}$ 0.665 0.638 0.650 0.645 0.588 0.620 0.633 0.525 $sR_{\text{BLEURT}}$ 0.647 0.602 0.626 0.607 0.615 0.608 0.604 0.500 $sF1_{\text{BLEURT}}$ 0.656 0.619 0.637 0.626 0.601 0.614 0.618 0.512 $sP_{\text{BERT}}$ 0.823 0.825 0.842 0.838 0.748 0.819 0.932 0.767 $sR_{\text{BERT}}$ 0.795 0.799 0.804 0.803 0.791 0.815 0.910 0.797 $sF1_{\text{BERT}}$ 0.809 0.811 0.823 0.820 0.769 0.818 0.921 0.781 Gemma FlanT5 LLaMA 3.1 (FT) LLaDA (FT) Mistral (FT) Gemma (FT) FlanT5 (FT) Arg-LLaDA $sP_{\text{BLEURT}}$ 0.618 0.415 0.640 0.643 0.545 0.652 0.545 0.671 $sR_{\text{BLEURT}}$ 0.642 0.390 0.631 0.616 0.530 0.637 0.559 0.645 $sF1_{\text{BLEURT}}$ 0.629 0.402 0.635 0.629 0.537 0.643 0.551 0.658 $sP_{\text{BERT}}$ 0.823 0.760 0.835 0.834 0.810 0.841 0.805 0.854 $sR_{\text{BERT}}$ 0.817 0.748 0.808 0.794 0.764 0.820 0.790 0.815 $sF1_{\text{BERT}}$ 0.822 0.754 0.822 0.813 0.784 0.830 0.797 0.834
Table 5: Detailed Soft-Score Evaluation on ArgKP. We report Soft-Precision, Soft-Recall, and Soft-F1 using two semantic similarity functions (BLEURT and BERTScore).